# Understanding In-context Learning of Addition via Activation Subspaces

### v2

@@ 7 Discussion
In this work, we provided a detailed mechanistic analysis of in-context learning for addition tasks in Llama-3-8B. We found that a small number of attention heads operating in low-dimensional subspaces can extract, represent, and aggregate ICL task information in structured and interpretable ways. We performed our analysis in five steps:
@@
    Examine this subspace qualitatively, which uncovered periodic patterns in the activation space (§4.2) that decomposed into interpretable subspaces encoding unit-digit and magnitude information (§4.3).
@@
There are a few other questions stemming from our work. First, the three important heads are capable of individually representing and extracting information, but Llama3-8B relies on all of them—how does it jointly aggregate information from these three heads to produce a final answer? Second, we observe that heads implement a “self-correction” mechanism indicated by the negative correlation between examples signals, but the exact mechanism of this self-correction remains to be discovered.
Our findings demonstrate how a focused study can reveal sophisticated latent structures and delicate computational strategies within a LLM. We hope this motivates more work to dive deep into the LLM inner structures and obtain more understanding of LLM latent spaces and computations.
@@ Acknowledgments
We thank Lisa Dunlap, Alex Pan, Tony Lian, Anya Ji, Jiaxin Ge, Junyi Zhang, Hanlin Zhu and Shu Liu for their helpful feedback and suggestions. KY is supported by the Vitalik Buterin Ph.D. Fellowship in AI Existential Safety. MJ is supported by the European Union (ERC-2022-SYG-OCEAN-101071601). LC is supported by a Miller Research Fellowship. JS is supported by Open Philanthropy, Simons Foundation, and NSF.
Appendix A Additional table in §3.2
In §3.2, we did systematic ablation studies to narrow down the tasks to three main heads from $33$ significant heads. The first step of the ablation studies is layer-wise ablation, where we mean-ablate the significant heads in a subset of layers. We include the experimental results here, which narrow down to layer 13 and 15.
@@
Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers. We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.87$, while other combinations lead to substantial drops in accuracy, especially when mean-ablating layers $13$ and $15$ (in blue).
Appendix B Additional figures in §4
In §4.2 and §4.3, we focused on head 1 for simplicity of exposition. Here we include the analogous experimental results for head 2 and 3.
B.1 Additional figures for §4.2
In §4.2, we found six trigonometric functions that can be linearly fitted by the coordinate functions of head 1 (Figure 6). Here we show analogous experimental results for head 2 and 3.
Figure 10: The coordinate functions (inner product with feature directions) of head 2’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
Figure 11: The coordinate functions (inner product with feature directions) of head 3’s vectors can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ well by linear transformation.
B.2 Additional figures for §4.3
In §4.3, we showed experimental results validating the hypothesis (ii) for head 1 (Figure 7). Here we show analogous experimental results validating the hypotheses (i) and (iii) for head 1, as well as all three hypotheses for heads 2 and 3. Recall that the three hypotheses are as follows:
@@ Appendix B Additional figures in §4
    the subspace spanned by by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
@@
We first show the experimental results validating hypotheses (i) and (iii) for head 1. Projecting out of the parity direction doesn’t lead to high errors for the parity and the final answer across all tasks, which might be because parity is relatively easy to obtain (e.g., random choice leads to 0.5 accuracy).
(a) Head 1 projected onto parity direction
(b) Head 1 projected out of parity direction
Figure 12: Validation of hypothesis (i) for head 1.
(a) Head 1 projected onto magnitude subspace
(b) Head 1 projected out of magnitude subspace
Figure 13: Validation of hypothesis (iii) for head 1.
We then show analogous results for head 2.
(a) Head 2 projected onto parity direction
(b) Head 2 projected out of parity direction
Figure 14: Validation of hypothesis (i) for head 2.
(a) Head 2 projected onto unit subspace
(b) Head 2 projected out of unit subspace
Figure 15: Validation of hypothesis (ii) for head 2.
(a) Head 2 projected onto magnitude subspace
(b) Head 2 projected out of magnitude subspace
Figure 16: Validation of hypothesis (iii) for head 2.
Finally, we show analogous results for head 3. The evidence for head 3 is slightly weaker than the first two heads, which aligns with the fact that the intervened accuracy of head 3 is slightly weaker than the ones of the first two heads (Tabel 1).
(a) Head 3 projected onto parity direction
(b) Head 3 projected out of parity direction
Figure 17: Validation of hypothesis (i) for head 3.
(a) Head 3 projected onto unit subspace
(b) Head 3 projected out of unit subspace
Figure 18: Validation of hypothesis (ii) for head 3.
(a) Head 3 projected onto magnitude subspace
(b) Head 3 projected out of magnitude subspace
Figure 19: Validation of hypothesis (iii) for head 3.

### v3

@@ 7 Discussion
In this work, we provided a detailed mechanistic analysis of in-context learning for addition tasks. We found that a small number of attention heads operating in low-dimensional subspaces can extract, represent, and aggregate ICL task information in structured and interpretable ways. We performed our analysis in five steps:
@@
    Examine subspace qualitatively, which uncovered periodic patterns in activation space (§4.2) that decomposed into interpretable subspaces encoding unit-digit and magnitude information (§4.3).
@@ Acknowledgments
We thank Lisa Dunlap, Alex Pan, Tony Lian, Ryan Wang, Jiaxin Ge, Anya Ji, Junyi Zhang, Hanlin Zhu, Shu Liu and Baifeng Shi for their helpful feedback and suggestions. KY is supported by the Vitalik Buterin Ph.D. Fellowship in AI Existential Safety. MJ is supported by the European Union (ERC-2022-SYG-OCEAN-101071601). LC is supported by a Miller Research Fellowship. JS is supported by Open Philanthropy, Simons Foundation, and NSF.
Appendix A Supplement for §3.1
A.1 Additional figure for §3.1
In §3.1, we developed a global optimization method to identify $33$ significant heads for Llama-3-8B-instruct. Here we visualize our optimized coefficients as well as the previous method, also comparing their intervention accuracy in Figure 3.
(a) Coefficients of heads.
(b) Average indirect effects of heads (Todd et al., 2024).
Figure 3: Comparison of significant heads identified by our optimized coefficients (left) and by average indirect effects (AIE) from the previous method (Todd et al., 2024) (right). Colors indicate the magnitude of each head’s importance (coefficients or AIE) on Llama-3-8B-instruct. The top $33$ heads identified by both methods are highlighted with frames ($13$ heads common across both methods in red and other $20$ heads in blue). Our identified heads yield an intervention accuracy of $0.85$, compared to the previous method’s accuracy of $0.31$. Both methods select heads from similar layers, but our optimization approach is significantly more effective.
A.2 Additional models for §3.1
We also train coefficients to get sets of important heads responsible for add-k on Llama-3.2-3B-instruct, Llama-3.2-3B, and Qwen-2.5-7B. In all cases, we achieve better accuracies with fewer heads than Todd et al. (2024). We report our results in Table 2.
Model Llama-3.2-3B-instruct Llama-3.2-3B Qwen-2.5-7B
ICL accuracy 0.50 0.64 0.91
Our weighted-heads accuracy 0.66 0.83 0.61
Our top-heads accuracy 0.62 0.75 0.34
Todd et al. (2024)’s top-heads accuracy 0.20 0.1 0.12
Table 2: Accuracies comparison between our method and Todd et al. (2024)’s method as well as the baseline ICL accuracy on add-k task cross other models. Our weighted-heads accuracy is the intervention accuracy achieved by the weighted sum of all heads where the weights are the raw coefficients at the last epoch of our training; our top-heads accuracy is the intervention accuracy achieved by the sum of top heads selected by coefficients at the last epoch of our training; and the Todd et al. (2024)’s top-heads accuracy is the intervention accuracy achieved by the sum of top heads selected by their average indirect effect. Here we choose the number of top heads as the one giving the highest accuracy for each case. For Llama-3 models, we need to choose around 30 heads while Todd et al. (2024)’s method needs to choose around 60 heads; for Qwen-2.5 model, we both need to choose $210$ heads. Our weighted-heads accuracy get significantly higher accuracy than Todd et al. (2024)’s, which indicates our method can find better function vectors than Todd et al. (2024)’s. Our top-heads accuracy are also higher with fewer heads than Todd et al. (2024)’s, which indicates our method can also find more effective set of heads for ICL tasks.
Appendix B Supplement for §3.2
B.1 Additional table for §3.2
In §3.2, we did systematic ablation studies to narrow down the tasks to three main heads from $33$ significant heads. The first step of the ablation studies is layer-wise ablation, where we mean-ablate the significant heads in a subset of layers. We include the experimental results here, which narrow down to layer 13 and 15 in Table 3.
@@
Table 3: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers on Llama-3-8B-instruct. We first narrow down to the layers before layer $15$, then the range of $[13,15]$ and finally $\{13,15\}$ (in red), which all almost preserve the clean accuracy of $0.87$, while other combinations lead to substantial drops in accuracy, especially when mean-ablating layers $13$ and $15$ (in blue).
B.2 Additional models for §3.2
We do the same ablation experiments for other three models, narrowing down to one layer and three heads for all of them. For two Llama-3.2 models, we narrow down to layer 14 and heads (14, 1), (14, 2), and (14, 12), the sum of which achieves accuracy $0.60$ and $0.70$. For Qwen-2.5-7B model, we narrow down to layer 21 and heads (21, 0), (21, 2), and (21, 5), the sum of which achieves accuracy $0.29$. Note that all of accuracies of these three heads are higher than the corresponding accuracies by Todd et al. (2024). We report the specific accuracies under different ablation setups in Tables 4, 5, 6, and 7.
Layer Llama-3.2-3B-instruct Accuracy Llama-3.2-3B Accuracy
No intervention 0.50 0.64
$[0,27]$ 0.62 0.75
$\{14\}$ 0.60 0.70
Other single layer (max) 0.06 0.05
Table 4: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers on Llama-3.2-3B-instruct and Llama-3.2-3B models. Both models narrow down to the layer $14$, which achieves significantly higher accuracy than any other single layer.
Head Coefficients Llama-3.2-3B-instruct Accuracy Llama-3.2-3B Accuracy
$(14,1)$ 5 / 5 0.78 0.95
$(14,2)$ 6 / 4 0.51 0.61
$(14,12)$ 4 / 4 0.26 0.26
Other single head Optimal 0.03 0.08
Table 5: Intervention accuracies for keeping the selected heads scaled by the corresponding coefficients and mean-ablating all the other significant heads on Llama-3.2-3B-instruct and Llama-3.2-3B models. We narrow down them both to the same three heads, which achieve significantly higher accuracy than any other single head in layer $14$.
Layer Qwen-2.5-7B Accuracy
No intervention 0.91
$[0,27]$ 0.34
$\{21\}$ 0.29
Other single layer (max) 0.03
Table 6: Intervention accuracies for keeping the significant heads in the selected layers and mean-ablating the significant heads in the remaining layers on Llama-3.2-3B-instruct and Llama-3.2-3B models. Both models narrow down to the layer $14$, which achieves significantly higher accuracy than any other single layer.
Head Coefficient Qwen-2.5-7B Accuracy
$(21,5)$ 4 0.29
$(21,0)$ 3 0.18
$(21,2)$ 4 0.15
Other single head Optimal 0.06
Table 7: Intervention accuracies for keeping the selected heads scaled by the corresponding coefficients and mean-ablating all the other significant heads on Qwen-2.5-7B. We narrow down to three heads, which achieve significantly higher accuracy than any other single head in layer 2121.
Appendix C Supplement for §4.1
We perform PCA on the $30$ task vectors and find that just six directions can explain $97\%$ of the task variance on Llama-3-8B-instruct (Figure 4(a)), and similarly for the three models (Figures 4(b), 4(c) and 4(d)).
(a) Llama-3-8B-instruct
(b) Llama-3.2-3B-instruct
(c) Llama-3.2-3B
(d) Qwen-2.5-7B
Figure 4: Explained variance ratio vs. number of PCs for each head across models. The first six PCs make up most of the explained variance (97%) for Llama-3 models, and the first eight PCs do so for Qwen-2.5-7B.
Appendix D Supplement for §4.2
We first present additional figures for Llama-3-8B-Instruct, followed by results for the other three models. The main finding is that three heads in each Llama-3.2 model and one head in the Qwen-2.5 model consistently encode periodic patterns through trigonometric functions. By contrast, the remaining two heads in Qwen-2.5 do not encode periodicity; instead, their coordinate functions exhibit distinct behaviors across the intervals $[1,10]$, $[11,20]$, and $[21,30]$. We conjecture that this difference arises from tokenization: Qwen encodes numbers digit by digit, leading to discontinuities across each 10-interval, whereas the Llama family represents every number below 100 as a single token.
D.1 Additional figures for §4.2
In §4.2, we found six trigonometric functions that can be linearly fitted by the coordinate functions. Here we first supplement the plot from which we observe the periodic pattern of the coordinate functions for the three heads (Figure 5) and then show the fitting functions for three heads (Figure 6).
(a) Head (15, 2)
(b) Head (15, 1)
(c) Head (13, 6)
Figure 5: Coordinates of three heads’ vectors (inner products with PCs) for the first six PCs across different add-$k$ tasks on Llama-3-8B-instruct. Periodic patterns are visible in the first few PCs of each head.
(a) Head (15, 2)
(b) Head (15, 1)
(c) Head (13, 6)
Figure 6: Coordinate functions of three heads can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ on Llama-3-8B-instruct.
D.2 Llama-3.2-3B-instruct for §4.2
The Llama family of models show similar results. We show results for Llama-3.2-3B-instruct here and omit Llama-3.2-3B.
(a) Head (14, 1)
(b) Head (14, 2)
(c) Head (14, 12)
Figure 7: Coordinates of three heads’ vectors (inner products with PCs) for the first six PCs across different add-$k$ tasks on Llama-3.2-3B-instruct. The first six PCs reveal clear periodic patterns.
(a) Head (14, 1)
(b) Head (14, 2)
(c) Head (14, 12)
Figure 8: Coordinate functions of three heads can fit six trigonometric functions with periods $2,5,10,10,25$, and $50$ on Llama-3.2-3B-instruct.
D.3 Qwen-2.5-7B for §4.2
One head (21, 0) shows similar periodic patterns while the other two heads do not encode periodicity; instead, their coordinate functions exhibit distinct behaviors across the intervals $[1,10]$, $[11,20]$, and $[21,30]$. We conjecture that this difference arises from tokenization: Qwen encodes numbers digit by digit, leading to discontinuities across each 10-interval, whereas the Llama family represents every number below 100 as a single token.
(a) Head (21, 0)
(b) Head (21, 2)
(c) Head (21, 5)
Figure 9: Coordinates of three heads’ vectors (inner products with PCs) for the first six PCs across different add-$k$ tasks on Qwen-2.5-7B. Head (21, 0) reveals periodic patterns, whereas Heads (21, 2) and (21, 5) show discontinuous behaviors across the intervals $[1,10]$, $[11,20]$, and $[21,30]$.
(a) Head (21, 0)
(b) Head (21, 2)
(c) Head (21, 5)
Figure 10: Six trigonometric functions with periods $2,5,10,10,25$, and $50$ fitted by coordinate functions of three heads on Qwen-2.5-7B. Only head (21, 0) fits the periodic functions well, while heads (21, 2) and (21, 5) do not, consistent with their non-periodic coordinate behaviors.
Appendix E Supplement for §4.3
Recall that the three hypotheses in §4.3 are as follows:
@@
    the subspace spanned by the directions with periods $25,50$, which we call the “magnitude subspace”, encodes the coarse magnitude (i.e., the tens digit) of $k$.
@@
We first show the experimental results validating hypothesis (ii) for head 1 (Figure 11), and then show analogous results for hypotheses (i) and (iii). Projecting out of the parity direction does not substantially increase errors for either the parity or the final answer across all tasks, which might be because parity is relatively easy to obtain (e.g., random choice leads to 0.5 accuracy).
Figure 11: The error rates for the unit digit and the final answer across tasks when projecting head 1’s vectors onto (left) and out of (right) the “unit subspace”. Projecting onto the unit subspace results in a low unit-digit error rate even when the final-answer error remains high, while projecting out leads to high unit-digit error rates that almost fully account for the final-answer errors. This confirms that the unit subspace specifically encodes the unit-digit signal.
(a) onto parity direction
(b) out of parity direction
(c) onto magnitude subspace
(d) out of magnitude subspace
Figure 12: Validation of hypotheses (i) and (iii) for head 1. Each row shows results for the parity and magnitude subspaces (left: projection onto; right: projection out of). The unit-digit hypothesis (ii) is omitted here since it is already shown in Figure 11.
(a) Head 2 onto parity direction
(b) Head 2 out of parity direction
(c) Head 2 onto unit subspace
(d) Head 2 out of unit subspace
(e) Head 2 onto magnitude subspace
(f) Head 2 out of magnitude subspace
Figure 13: Validation of hypotheses (i)–(iii) for head 2. Each row shows the projection effects for the parity, unit, and magnitude subspaces respectively.
(a) Head 3 onto parity direction
(b) Head 3 out of parity direction
(c) Head 3 onto unit subspace
(d) Head 3 out of unit subspace
(e) Head 3 onto magnitude subspace
(f) Head 3 out of magnitude subspace
Figure 14: Validation of hypotheses (i)–(iii) for head 3. The evidence is slightly weaker than for heads 1 and 2, consistent with head 3’s lower intervened accuracy (Table 1).
Appendix F Supplement for §5.2
We show the strength and direction of the signals extracted from each individual token through a random mixed-k ICL prompt as an example below. Since we find out the Llama family of models behave similarly in the previous sections, we here just show results for Llama-3-8B-instruct and Qwen-2.5-7B as examples.
F.1 Additional figures for §5.2
All three heads for Llama-3-8B-instruct behave similarly in the signal strengths and directions. They all peak at the $y$ tokens and extract signal corresponding to $y_{i}-x_{i}$.
(a) Head (15, 2)
(b) Head (15, 1)
(c) Head (13, 6)
Figure 15: Llama-3-8B-instruct. For each head, the top panel shows the strength of task-signal contribution of each previous token $t$ to the final token, $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$. Decomposing it into two parts: (1) the norm of the extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$ and (2) the attention score from the final token $\alpha_{t}$, both consistently peak at the tokens $t=y_{i}$ (in bold). The bottom panel shows the inner product between the projected signal from each $y_{i}$ and the head vector $\tilde{h}_{k}$ (normalized to unit norm), which peaks at $k=y_{i}-x_{i}$, indicating that each head extracts $y_{i}-x_{i}$ from its corresponding demonstration $x_{i}\!\to\!y_{i}$.
F.2 Qwen-2.5-7B for §5.2
Notice that Qwen-2.5-7B has a different tokenizer from the Llama models: it tokenizes numbers digit by digit. The three heads behave qualitatively similarly—peaking at $y$ tokens and extracting $y_{i}-x_{i}$—though sometimes with an offset of $10$ or $20$.
(a) Head (21, 0)
(b) Head (21, 2)
(c) Head (21, 5)
Figure 16: Qwen-2.5-7B. For each head, the top panel shows $\|\alpha_{t}W_{h}O_{h}V_{h}z_{t}\|$, decomposed into two parts: (1) the norm of extracted information $\|W_{h}O_{h}V_{h}z_{t}\|$, and (2) the attention score from the final token $\alpha_{t}$. Both consistently peak at the tokens $t=y_{i}$ (in bold). The bottom panel shows the inner products between the projected signals from $y_{i}$ and head vectors $\tilde{h}_{k}$, which peak near $k=y_{i}-x_{i}$ (sometimes offset by 10 or 20), consistent with Qwen’s digit-wise tokenization.
Appendix G Supplement for §5.3
We report the concrete numeric variables computed in §5.3 for Llama-3-8B-instruct, Llama-3.2-3B, and Qwen-2.5-7B. The self-correction mechanism significantly exists in Llama family of models while partially exists in Qwen model.
G.1 Additional table for §5.3
The signals from any two demonstrations are mostly negatively correlated for all three heads, suggesting a self-correction mechanism on Llama-3-8B-instruct.
Stat Neg Pos
Avg -2.01 0.27
Min -1.40 0.07
Max -2.34 0.54
(a) Head (15, 2)
Stat Neg Pos
Avg -1.95 0.05
Min -1.65 0.00
Max -2.17 0.20
(b) Head (15, 1)
Stat Neg Pos
Avg -0.76 0.28
Min -0.15 0.00
Max -1.68 1.03
(c) Head (13, 6)
Table 8: Llama-3-8B-instruct: Statistics (min, max, average) of the absolute values of negative and positive correlation sums over the 30 tasks for three heads. The negative correlation sum is significantly higher for all heads, indicating that signals from any two demonstrations are mostly negatively correlated — a hallmark of the self-correction mechanism.
G.2 Llama-3.2-3B for §5.3
The signals from any two demonstrations are mostly negatively correlated for the two stronger heads, consistent with the self-correction mechanism observed in Llama-3-8B-instruct.
Stat Neg Pos
Avg -2.12 0.23
Min -1.96 0.00
Max -2.41 0.53
(a) Head (14, 1)
Stat Neg Pos
Avg -2.07 0.05
Min -1.89 0.00
Max -2.24 0.22
(b) Head (14, 2)
Stat Neg Pos
Avg 0.27 0.75
Min 0.00 0.00
Max -0.90 0.75
(c) Head (14, 12)
Table 9: Llama-3.2-3B: Statistics (min, max, average) of the absolute values of negative and positive correlation sums over the 30 tasks for three heads. Heads (14, 1) and (14, 2) show clear negative correlation dominance, while head (14, 12) does not — consistent with its weaker task accuracy (Table 5).
G.3 Qwen-2.5-7B for §5.3
For Qwen-2.5-7B, the correlation pattern varies by head. The head encoding periodic patterns (21, 0) still shows strong negative correlations, while the others exhibit weaker or even positive correlations—indicating that the self-correction mechanism only partially exists.
Stat Neg Pos
Avg -1.84 0.08
Min -1.44 0.00
Max -2.15 0.67
(a) Head (21, 0)
Stat Neg Pos
Avg -0.64 0.50
Min -1.44 0.03
Max -2.19 1.57
(b) Head (21, 5)
Stat Neg Pos
Avg -0.37 2.62
Min 0.00 0.00
Max -1.85 5.54
(c) Head (21, 2)
Table 10: Qwen-2.5-7B: Statistics (min, max, average) of the absolute values of negative and positive correlation sums over the 30 tasks for three heads. Head (21, 0) shows clear negative correlations, while Heads (21, 5) and (21, 2) show mixed or positive correlations, suggesting that self-correction partially exists in Qwen-2.5-7B.